OpenStudy (anonymous):

will fan and medal Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10, and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (anonymous):

It gives you an equation to use with all the needed information to solve it. a= adult c=child So basically all you have to do is substitute C with the 5 children presented in the question. 10A+6(5)= 80 (SUBSTITUTE) 10A+30= 80 -30 -30 (SUBTRACT) 10A= 50 --- --- (DIVIDE) 10 10 A= 5 Sooooo 5 adult tickets

OpenStudy (anonymous):

tadaaaaaaa